OpenStudy (anonymous):

find the xintercepts and vertext of the porabola y=x^2+2x-24

OpenStudy (anonymous):

y=x^2+2x-24 y=(x-4)(x+6) for x intercepts, y=0, 0=(x-4)(x+6) x1=4 and x2=-6 ans

OpenStudy (anonymous):

from y=x^2+2x-24, a=1,b=2,c=-24 vertex v= -b/2a=-2/(2)(1)=-1